OpenStudy (anonymous):

Solve for X: 4x-7/3=x-9

OpenStudy (alexandervonhumboldt2):

first 4x-7/3=x-9 subtartc x from both sides. you get 3x-7/3=-9 now add 7/3 to both sides: 3x=-27/3+7/3 3x=-20/3 now divide by 3 both sides: x=-20/9

OpenStudy (anonymous):

\[4x-\frac{ 7 }{ 3 }=x-9\]

OpenStudy (anonymous):

Yay ok thnks so much :) @AlexandervonHumboldt2 @DaWizjr

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (alexandervonhumboldt2):

☺☻

OpenStudy (anonymous):

yea